Circular No. 5814

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA IN M31 R. R. Treffers, B. Leibundgut, and A. V. Filippenko, University of

California at Berkeley; and M. W. Richmond, Princeton University, report the discovery of a probable nova in M31 at α = 0^{h}40^{m}00^{s}, δ = +40^{o}59′.7 (equinox 1950.0), which is ∼ 55′′ east and 79′′ north of the galaxy’s nucleus. The object was found during the Leuschner Observatory Supernova Search, which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera, and it was detected in an image obtained on June 9 UT at magnitude R = 15.8 ± 0.3. Confirmation was obtained on June 11. No object appears at this position to a limit of mag ∼ 18 in an exposure taken on May 29.

V1974 CYGNI F. Paresce, Space Telescope Science Institute and European Space

Agency, and the Faint Object Camera Investigation Definition Team communicate: “The Faint Object Camera on the Hubble Space Telescope observed V1974 Cyg (Nova Cyg 1992) on May 31 in the ultraviolet (FOC filter F278M, whose bandpass peaks at 0.28 nm and encompasses several emission lines of ionized neon, magnesium, and oxygen between 0.25 and 0.3 nm). The image reveals a circular ring of emission of radius 0′′.130 ± 0′′.020 centered on the star. The azimuthal surface brightness profile peaks in the northwest sector of the ring at a level comparable to that of the central star and a factor of two above that of the southeast sector at these wavelengths. A bridge of emission links the star to the ring at position angle 30^{o}. Residual emission in the region between the ring and the star that is not due to spillover from the wings of the point spread function indicates that the ring is most likely due to a thin, inhomogeneous, limb-brightened shell of emitting material. A preliminary estimate of the brightness ratio of the limb and the half-radius point yields an upper limit to the shell thickness of 10 percent of the radius. Comparison with observations taken soon after outburst (IAUC 5463) show that the shell has been expanding at a constant angular rate of 0′′.00028/day. Assuming an average velocity of expansion of 1500 km/s, the observed ring location and epoch imply a source distance of 3.2 kpc.”

1993 June 11 Daniel W. E. Green